OpenStudy (anonymous):

simplify sec^2(theta)-cos^2(theta)*sec^2(theta)

OpenStudy (bahrom7893):

Keep in mind that Sec(theta) = 1/Cos(theta)

OpenStudy (bahrom7893):

Which means that: Sec^2(theta) = 1^2/Cos^2(theta) = 1/Cos^2(theta)

OpenStudy (bahrom7893):

Plugging that into your equation: 1/Cos^2(theta) - Cos^2(theta) * 1/Cos^2(theta) = 1/Cos^2(theta) - 1

OpenStudy (bahrom7893):

If we simplify this a bit further by finding the common denominator, we get: [1-Cos^2(theta)]/Cos^2(theta) = Sin^2(theta)/Cos^2(theta) = Tan^2(theta)

OpenStudy (bahrom7893):

Since Sin^2(theta)+Cos^2(theta)=1

OpenStudy (anonymous):

thankyou